OpenStudy (anonymous):

show that with respect to point wise addition and multiplication of function ,the set of all differentiable function f:R TO R is a commutative ring with unit element?

terenzreignz (terenzreignz):

So tell me... what exactly IS a commutative ring? :)

OpenStudy (anonymous):

a commutative ring with unit element which has at least two elements ,but nozero divisor is called integral domain

terenzreignz (terenzreignz):

You're back. We don't need an integral domain, we just need a commutative ring, which, by definition, is a ring with its multiplication element also commutative. Okay?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Now... what are the two operations defined on functions in this case?

OpenStudy (anonymous):

addition and multiplication

terenzreignz (terenzreignz):

Yes. IS Addition commutative on functions?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

What about multiplication?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Let f and g be differentiable functions from R to R. Is it not so that \[\Large f(x)g(x) = g(x)f(x)\]?

OpenStudy (anonymous):

yes but in ring does not equal to each other

terenzreignz (terenzreignz):

What do you mean?

OpenStudy (anonymous):

i mean the ring have to has all the properties except mult identity ,mult commutative and mult inverse

terenzreignz (terenzreignz):

How is it not commutative under multiplication? Can you give an example of two differentiable functoins from \(\large \mathbb{R}\rightarrow \mathbb{R}\) such that \[\Large f(x)g(x) \ne g(x)f(x)\]?

OpenStudy (anonymous):

yes such that if i have matrices 1 2 3 0 and 1 3 2 4 doesn't equal to 1 3 2 4 times 1 2 3 0

terenzreignz (terenzreignz):

But we aren't dealing with matrices are we? We're dealing with DIFFERENTIABLE FUNCTIONS from \(\large \mathbb{R}\rightarrow \mathbb{R}\)

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

Granted that multiplication of functions is COMMUTATIVE. Do you agree?

OpenStudy (anonymous):

yes i am confused because the ring have to don't has multipcation commutative and in this question i have toprove that is commautive ring with unit element

terenzreignz (terenzreignz):

A ring doesn't HAVE TO BE a commutative ring (IE, its 'multiplication' operator doesn't have to be commutative). But it CAN be. Do you understand?

OpenStudy (anonymous):

can you give me example of set and prove that to be more clear

terenzreignz (terenzreignz):

The set of Integers form a commutative ring.

terenzreignz (terenzreignz):

That's because (1) When you add any two integers, the sum is an integer (2) When you multiply any two integers, the product is an integer (3) There is an integer, call it 0, such that when you add it to any integer, the sum is THAT integer (4) For any integer a, there is an integer, call it -a such that a + (-a) = 0 (5) For any 2 integers a, b, a+b = b+a (6) For any 2 integers, a, b, ab = ba (7) For any 3 integers, a, b, and c, a(b + c) = ab + ac The conclusion holds.

OpenStudy (anonymous):

okay if i choose two element like give me any two elements in intger because when i choose 1 and 2 it does not have addition closure and i do not know how i choose intger or when i want to choose z mode 3 how i can work with mult idinity in this case

terenzreignz (terenzreignz):

^I'm sorry, but I really did not understand that...

OpenStudy (anonymous):

sorry i am international student and maybe because of that.. i meant my problem can you give me number in integer i can work with because i had work with z mod 3 and when come to deal with mult identity i do not know how .. hope its clear

terenzreignz (terenzreignz):

You need to work with this group/ring? \[\Large \mathbb{Z}_3\]

OpenStudy (anonymous):

okay can you explain to me mult identity in this group

terenzreignz (terenzreignz):

It's 1. And anyway, that's beside the point... we still have not yet finished with your original question.

OpenStudy (anonymous):

yes you mean i have to choose any number in the set z mode 3 and multiply in 1 ?

terenzreignz (terenzreignz):

We should save this for later. First, we work on your original question, because this one is rather easy compared to the original. Is function multiplication distributive over addition? In other words, for any functions f,g, and h, is it true that \[\Large f(x)[g(x) + h(x) ]= f(x)g(x) +f(x)h(x)\]?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay, two more things. When you add two differentiable functions, is the sum also differentiable?

OpenStudy (anonymous):

yes sure

terenzreignz (terenzreignz):

(Note, we're skipping through details, I'm assuming you can fill them in) Finally, when you multiply two differentiable functions, is the product differentiable?

OpenStudy (anonymous):

not really

terenzreignz (terenzreignz):

Counterexample?

OpenStudy (anonymous):

excuse me you do not understand where i am struggle with i man can you choose any two element number and show me how you can prove dors not have mult idinity ,mult commautivity and mult invers and in the same time it is ring commuativity with unit elements

terenzreignz (terenzreignz):

Calm yourself, and take it slow. You said the product of two differentiable functions from \(\large \mathbb{R}\rightarrow \mathbb{R}\) is not necessarily differentiable. Then you can provide an example?

OpenStudy (anonymous):

i have to choose two different elements to prove that

terenzreignz (terenzreignz):

Yes... can you?

terenzreignz (terenzreignz):

Can you find two differentiable functions whose product is not differentiable?

OpenStudy (anonymous):

okay if i choose from z mode 3 that 2 mode 3 + 1 mode 3 =0 and its belong to z mode 3

terenzreignz (terenzreignz):

Please don't change the subject. We are talking about differentiable functions from \(\large \mathbb{R}\rightarrow \mathbb{R}\) . And me asking you to find two such functions whose product is NOT differentiable. Please stop bringing up \(\large \mathbb{Z}_3\), for the moment, it is NOT RELEVANT to the discussion.

OpenStudy (anonymous):

1+2=3

terenzreignz (terenzreignz):

^Please read my previous post.

OpenStudy (anonymous):

can i choose matrices

terenzreignz (terenzreignz):

ARE matrices differentiable functions? NO! They are an entirely different notion.

OpenStudy (anonymous):

sorry about that can you give me that...confused

terenzreignz (terenzreignz):

Let's backtrack a bit. If f and g are two differentiable functions from \(\large \mathbb{R}\rightarrow \mathbb{R}\) is f(x)g(x) also differentiable?

terenzreignz (terenzreignz):

And where are you from, btw?

OpenStudy (anonymous):

Saudi Arabia first time study math in English

terenzreignz (terenzreignz):

Right. Anyway, my question before I asked you where you're from. Think about it.

OpenStudy (anonymous):

think about what??

OpenStudy (zzr0ck3r):

f,g differentiable then is f*g differentiable was the question...

terenzreignz (terenzreignz):

Let's backtrack a bit. If f and g are two differentiable functions from R→R is f(x)g(x) also differentiable?

terenzreignz (terenzreignz):

You might want to recall a nifty rule called 'the product rule' \[\Large \frac{d}{dx}f(x)g(x) = f(x)g'(x) + f'(x)g(x) \]

terenzreignz (terenzreignz):

(Even though you really shouldn't have to T.T)

OpenStudy (anonymous):

what is T.T

terenzreignz (terenzreignz):

An emoticon. Pay no heed to it, pay attention ONLY to my question.

OpenStudy (anonymous):

SO ANYWAY thank you for your cooperation

OpenStudy (anonymous):

i understand that it has to be differentiable and you do not answer my question can you explain to me how i can get two elements and prove that its ring with unite elements and commutative

terenzreignz (terenzreignz):

For which?

OpenStudy (anonymous):

The important thing I am trying to understand to prove any set that is commutative ring with unit elements

terenzreignz (terenzreignz):

Fair enough,but that's easily determined... what is the differentiable function e(x) such that e(x)f(x) = f(x) for all differentiable functions f?

OpenStudy (anonymous):

1 times 2 =2 like this

terenzreignz (terenzreignz):

Yes... so what is e(x)?

OpenStudy (anonymous):

1

terenzreignz (terenzreignz):

And that's the unity.

OpenStudy (anonymous):

okay that is what i want the rule for any set to be ring must does not have mult identity and does not have mult commutative can you prove that

OpenStudy (watchmath):

mshms, you need to read your book again and get better understanding of the definition

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i will read it a gain

terenzreignz (terenzreignz):

Please do :)

OpenStudy (anonymous):

thanks